Given a triangle with sides 7m, 8m and 3m. What is 50% the area of this triangle?


1
Expert's answer
2021-09-13T16:33:36-0400

Using Heron's formula ;

Area(Triangle) ="\sqrt {s(s-a) (s-b) (s-c)}"

Where "s=\frac{(7+8+3) }{2} =9"

Area Triangle ="\sqrt {9(9-7) (9-8) (9-3)}"

Area Triangle ="\sqrt {108}=10.39m^2"

50%Area Triangle ="0.5\u00d710. 39m^2" "=5.196cm2"